OpenStudy (anonymous):

Can someone help? I'm a confused. If g(x) = x^2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

well just put the value in the function\[g(3)=3^2+2=?\]

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

I got g=11/3

OpenStudy (anonymous):

oh No, the problem do not asks you to evaluate g but g(3)

OpenStudy (anonymous):

here g(x) is a function and x is independent variable

OpenStudy (anonymous):

use 3 in place of x and evaluate (solve)

OpenStudy (phi):

algebra is confusing sometimes. when you see g(x) = x^2 +2 the g(x) does not mean g times x so you have to remember: g(x) (written exactly that way) means g(x) is a short-cut name for x^2 +2 when they write g(3), that is a short-cut way to write 3^2 +2 (they replace the x with a 3) the answer is just the part: 3^2+2 = 11

OpenStudy (anonymous):

Wow, thank you guys! @mukushla thank you for helping me so I know what to do. And thank you @phi for going step by step!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

So then, heres another, to see if I can get it. The question is if g(x) = 3(x - 2), find the value of x if g(x) = 6

OpenStudy (anonymous):

It would be 4? Right?

OpenStudy (phi):

g(x) = 3(x - 2) replace g(x) with 6 6= 3(x-2) divide both sides by 3 2 = x-2 add 2 to both sides 4= x yes, x is 4

OpenStudy (anonymous):

Awesome!! THANK YOU!